OpenStudy (anonymous):

The sum of three consecutive even integers is 312. Find the integers. Write the three integers in order of smallest to largest:

OpenStudy (wolf1728):

x + (x+1) + (x+2) = 312 3x + 3 = 312 3x = 309 x = 103 So the three integers are: 103 104 105

OpenStudy (anonymous):

but there even

geerky42 (geerky42):

"three consecutive even integers"

OpenStudy (anonymous):

those are odd

geerky42 (geerky42):

you can do similar thing to wolf's work though. x + (x+2) + (x+4) = 312

OpenStudy (anonymous):

\[x+(x+2)+(x+4) = 312\] Solve for x.

OpenStudy (anonymous):

156

OpenStudy (anonymous):

three consecutive

OpenStudy (anonymous):

cant ever be even

geerky42 (geerky42):

did you try and solve for x?

OpenStudy (anonymous):

yea hold on

OpenStudy (anonymous):

x3+3=312

OpenStudy (anonymous):

or is it x2+2=312

OpenStudy (anonymous):

idk

geerky42 (geerky42):

none. \(x + (x+2) + (x+4)=312 \\ x+x+x+2+4=312\\3x+6=312\)

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

306

OpenStudy (anonymous):

??

OpenStudy (anonymous):

102,104 106

geerky42 (geerky42):

you got it right

OpenStudy (anonymous):

or no im 2 off

OpenStudy (anonymous):

no nvm

geerky42 (geerky42):

102, 104, 106 are correct